# Output-Sensitive Evaluation of Regular Path Queries

### v1

@@ Abstract.
The Product Graph (PG) is the established evaluation approach for regular path queries. PG first constructs the product automaton of the data graph and the query and then uses breadth-first search to find the accepting states reachable from each starting state in the product automaton. Its data complexity is $O(|V|\cdot|E|)$, where $V$ and $E$ are the sets of nodes and respectively edges in the data graph. This complexity cannot be improved by combinatorial algorithms.
In this paper, we introduce OSPG, an output-sensitive refinement of PG, whose data complexity is $O(|E|^{3/2}+\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$, where OUT is the number of distinct node pairs in the query output. OSPG’s complexity is at most that of PG and can be asymptotically smaller for small output and sparse input. The improvement of OSPG over PG is due to the unnecessary time wasted by PG in the breadth-first search phase, in case a few output pairs are eventually discovered. For queries without Kleene star, the complexity of OSPG can be further improved to $O(|E|+|E|\cdot\sqrt{\textsf{OUT}})$.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 44th ACM SIGMOD-SIGACT-SIGAI Symposium on Principles of Database Systems; June 22-27, 2025; Berlin, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Graph-based database modelsccs: Theory of computation Database query processing and optimization (theory)ccs: Theory of computation Graph algorithms analysis
@@ 1. Introduction
Regular path queries are an essential construct in graph query languages (Angles et al., 2017; Deutsch et al., 2022), including industry efforts such as SPARQL from W3C (Harris and Seaborne, 2013), Cypher from Neo4J (Francis et al., 2018), GSQL from TigerGraph (TigerGraph Team, 2021), PGQL from Oracle (Property Graph Query Language, 2021), and more recently the standards SQL/PGQ and GQL (Deutsch et al., 2022; Francis et al., 2023). Such queries have also been investigated to a great extent in academia, e.g., (Angles et al., 2017; Baeza, 2013; Calvanese et al., 2003; Wood, 2012), since their introduction as a natural declarative formalism for expressing path matching in graphs (Mendelzon and Wood, 1989).
@@
The classical semantics can be realized by a simple and effective algorithm, called Product Graph (PG(Mendelzon and Wood, 1995; Baeza, 2013; Martens and Trautner, 2018). This is the workhorse of practical implementations in several existing graph database systems, e.g., MillenniumDB (Vrgoc et al., 2023), PathFinder (Calisto et al., 2024), TigerGraph (Deutsch et al., 2022), and RelationalAI. PG first constructs the product graph (product automaton) $P_{G,Q}$ of the data graph $G=(V,E,\Sigma)$ and the non-deterministic finite automaton $M_{Q}=(V_{Q},E_{Q})$ representing the query $Q$. This product graph has vertices $V\times V_{Q}$ and an edge $((v,q),(u,p))$ iff, for some symbol $\sigma\in\Sigma$, $G$ and $M_{Q}$ have $\sigma$-labeled edges $(v,\sigma,u)$ and respectively $(q,\sigma,p)$. PG then uses graph searching techniques like breadth-first search to find the accepting states reachable from each starting state in the product graph $P_{G,Q}$. The two steps of PG can be intertwined, so that the product graph need not be materialized.
@@
In this article, we introduce OSPG, a refinement of PG that is output-sensitive in the sense that its running time depends on the size of the query output. The data complexity of OSPG is $O(|E|^{3/2}+\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$, where OUT is the query output size. OSPG achieves this complexity without knowing a priori the size of the query output. Moreover, this complexity is at most that of PG and can be asymptotically smaller for sparse input and small output. This regime includes common scenarios, where the query is selective and the number of edges is far from maximum. In this regime, the improvement of OSPG over PG is due to the unnecessary time wasted by PG in the breadth-first search phase when only a few output pairs are eventually discovered.
@@ 2. Preliminaries
In the literature, given a regular path query $Q$ over an input graph $G=(V,E,\Sigma)$, it is standard to consider the product graph $P_{G,Q}$ of $G$ with an NFA $M_{Q}=(V_{Q},E_{Q},\Sigma)$ that recognizes the language $L(Q)$. The start states of $P_{G,Q}$ are pairs $(v,q)$ where $v$ is a vertex in $G$ and $q$ is a starting state of $M_{Q}$. The final (or accepting) states of $P_{G,Q}$ are pairs $(v,q)$ where $v$ is a vertex in $G$ and $q$ is a finish state of $M_{Q}$. The goal is to find paths from starting states to finish states in the product graph.
Definition 2.5 ($Q$-reachable and $Q$-degree of a vertex in an edge-labeled graph).
@@ 4.1. Reduction from any RPQ $Q$ to the RPQ $ab^{*}c$
Our reduction from any RPQ $Q$ to the RPQ $ab^{*}c$ essentially shows that $ab^{*}c$ is the hardest RPQ. Let $M_{Q}=(V_{Q},E_{Q},\Sigma)$ be an NFA for $Q$, represented as an edge-labeled graph (cf. Section 2). First, we construct the product graph $\overline{G}\stackrel{{\scriptstyle\text{def}}}{{=}}G\times M_{Q}$. Let $\overline{V}$ and $\overline{E}$ be the set of vertices and edges of $\overline{G}$, respectively. Then, we construct an edge-labeled graph $G^{\prime}=(V^{\prime},E^{\prime},\Sigma^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\{a,b,c\})$ as follows: The set of vertices $V^{\prime}$ is the same as the set of vertices $\overline{V}$. The set of edges $E^{\prime}$ contains the same edges in $\overline{E}$ but with the additional label ``$b"$. Moreover, $E^{\prime}$ contains a self-loop labeled ``$a"$ for each vertex corresponding to a start state in the NFA $M_{Q}$, and a self-loop labeled ``$c"$ for each vertex corresponding to a final state in $M_{Q}$. Formally:
@@
$\{((v,q),``a",(v,q))\mid\text{$v\in V$ and $q$ is a start state in the NFA $M_{Q}$}\}\quad\cup$
$\{((v,q),``c",(v,q))\mid\text{$v\in V$ and $q$ is a final state in the NFA $M_{Q}$}\}$
@@
Suppose we have an edge-labeled graph $G=(V,E,\Sigma=\{d,e,f,g\})$ and consider the RPQ $Q=d^{*}(e\cdot f+g)^{*}$. We demonstrate the above reduction from $Q$ to the RPQ $ab^{*}c$ over another edge-labeled graph $G^{\prime}=(V^{\prime},E^{\prime},\Sigma^{\prime}=\{a,b,c\})$. The regular expression $Q$ can be translated into the DFA $M_{Q}=(V_{Q},E_{Q},\Sigma)$ that is depicted in Figure 1. This DFA has states $\{q_{0},q_{1},q_{2}\}$, an initial state $q_{0}$, final states $\{q_{0},q_{2}\}$, and the following transitions, given as labeled edges:
@@ 4.2. Solving the RPQ $ab^{*}c$
Given a vertex $X$, we define $\deg_{R}(X)$ to be the number of distinct $Y$ values satisfying $R(X,Y)$. In the second step of the algorithm, we partition vertices $X\in V$ based on $\deg_{R}(X)$. In particular, a vertex $X$ is called light if $\deg_{R}(X)\leq\sqrt{|E|}$ and heavy otherwise. Note that by definition of $R$ and Definition 2.5, this is equivalent to saying that a vertex $X$ is light if its $(b^{*}c)$-degree is at most $\sqrt{|E|}$ and heavy otherwise. We define $R_{\ell}$ to be the set of pairs of vertices $(X,Y)\in V\times V$ where $X$ is light and $Y$ is $(b^{*}c)$-reachable from $X$, as shown in Eq. (3). We also define $R_{h}$ in Eq. (4) to be the set of heavy vertices $X\in V$. Given $R$, both $R_{\ell}$ and $R_{h}$ are straightforward to compute.
Finally, the algorithm evaluates the Datalog program given by Eq. (5)–(8), which defines two binary relations $Q_{\ell}$ and $Q_{h}$, and the algorithm returns their union as the final output to the RPQ. We show below that $Q_{\ell}$ contains all pairs of vertices $(v,u)$ where $u$ is $(ab^{*}c)$-reachable from $v$ through a path that only visits light vertices. In contrast, $Q_{h}$ contains all pairs of vertices $(v,u)$ where $u$ is $(ab^{*}c)$-reachable from $v$ through a path that visits at least one heavy vertex.
@@
4:Evaluate $R(X,Y)$ below while maintaining for every $X$, at most $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{|E|}+1$ distinct $Y$ values:
@@
7:Compute $Q_{\ell}$ and $Q_{h}$:
@@ 5. Complexity Analysis
Thanks to the reduction in Section 4.1, we only need to analyze the runtime of Algorithm 1 for the special RPQ $ab^{*}c$. We start with analyzing the runtime needed for the first step of Algorithm 1, which computes the relation $R(X,Y)$. Recall that for each vertex $X$, $R$ contains at most $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{|E|}+1$ different vertices $Y$ that are $(b^{*}c)$-reachable from $X$. Therefore, $|R|\leq|V|\cdot\Delta$. In order to compute $R$, we maintain, for every vertex $X$, a list $\mathsf{list}(X)$ of up to $\Delta$ different vertices $Y$ that are $(b^{*}c)$-reachable from $X$, as follows:
@@
The size of $S_{h}$ cannot exceed $\min\left(\frac{\textsf{OUT}}{\sqrt{|E|}},|V|\right)$. This is because for every vertex $X\in S_{h}$, there are more than $\sqrt{|E|}$ different vertices $Y$ such that the pair $(X,Y)$ is in the output of the RPQ $ab^{*}c$, and the total number of such pairs $(X,Y)$ is OUT by definition. By Rule (6) and then inductively by Rule (7), we have $\pi_{X}T(X,Y)\subseteq S_{h}$. By Rule (8), we have $\pi_{X}Q_{h}(X,Y)\subseteq\pi_{X}T(X,Z)\subseteq S_{h}$. For every fixed $X$-value in $S_{h}$, Rules (6), (7) and (8) take time $O(|E|)$ in total because they correspond to a single-source traversal of the input graph, where each edge will be traversed at most once. Hence, the total time over all $X$-values in $S_{h}$ is $O\left(\min\left(\frac{\textsf{OUT}}{\sqrt{|E|}},|V|\right)\cdot|E|\right)=O(\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$, as desired. ∎
6. When is OSPG better than PG?
The data complexity [^6] This also holds for the combined complexity. of OSPG is at most that of PG and can be asymptotically smaller for sparse input and small output: This holds for $\textsf{OUT}=O((|V|\cdot\sqrt{|E|})^{1-\alpha})$ and $|E|=O(|V|^{2-\beta})$ for any $\alpha\in(0,1]$ and $\beta\in(0,2)$. This regime includes common scenarios, where the query is selective and the number of edges is far from maximum. In this regime, the improvement of OSPG over PG is due to the unnecessary time used by PG in the breadth-first search phase when only a few output pairs are eventually discovered. We next exemplify this gap between OSPG and PG in two cases.

### v2

@@ Abstract.
The Product Graph (PG) is the established evaluation approach for regular path queries. PG first constructs the product automaton of the data graph and the query and then uses breadth-first search to find the accepting states reachable from each initial state in the product automaton. Its data complexity is $O(|V|\cdot|E|)$, where $V$ and $E$ are the sets of vertices and respectively edges in the data graph. This complexity cannot be improved by combinatorial algorithms.
In this paper, we introduce OSPG, an output-sensitive refinement of PG, whose data complexity is $O(|E|^{3/2}+\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$, where OUT is the number of distinct vertex pairs in the query output. OSPG’s complexity is at most that of PG and can be asymptotically smaller for small output and sparse input. The improvement of OSPG over PG is due to the unnecessary time wasted by PG in the breadth-first search phase, in case a few output pairs are eventually discovered. For queries without Kleene star, the complexity of OSPG can be further improved to $O(|E|+|E|\cdot\sqrt{\textsf{OUT}})$.
@@
ccs: Information systems Graph-based database modelsccs: Theory of computation Database query processing and optimization (theory)ccs: Theory of computation Graph algorithms analysis
@@ 1. Introduction
Regular path queries are an essential construct in graph query languages (Angles et al., 2017; Deutsch et al., 2022), including industry efforts such as SPARQL from W3C (Harris and Seaborne, 2013), Cypher from Neo4J (Francis et al., 2018), GSQL from TigerGraph (TigerGraph Team, 2021), PGQL from Oracle (Property Graph Query Language, 2021), and more recently the standards SQL/PGQ and GQL (Deutsch et al., 2022; Francis et al., 2023). Such queries have also been investigated to a great extent in academia, e.g., (Angles et al., 2017; Baeza, 2013; Calvanese et al., 2003; Wood, 2012), since their introduction as a natural declarative formalism for expressing path matching in graphs (Mendelzon and Wood, 1995).
@@
The classical semantics can be realized by a simple and effective algorithm, called Product Graph (PG(Mendelzon and Wood, 1995; Baeza, 2013; Martens and Trautner, 2018). This is the workhorse of practical implementations in several existing graph database systems, e.g., MillenniumDB (Vrgoc et al., 2023), PathFinder (Calisto et al., 2024), TigerGraph (Deutsch et al., 2022), and RelationalAI. PG first constructs the product graph (product automaton) $P_{G,Q}$ of the data graph $G=(V,E,\Sigma)$ and the non-deterministic finite automaton $M_{Q}=(V_{Q},E_{Q})$ representing the query $Q$. This product graph has vertices $V\times V_{Q}$ and an edge $((v,q),(u,p))$ iff, for some symbol $\sigma\in\Sigma$, $G$ and $M_{Q}$ have $\sigma$-labeled edges $(v,\sigma,u)$ and respectively $(q,\sigma,p)$. PG then uses graph searching techniques like breadth-first search to find the accepting states reachable from each initial state in the product graph $P_{G,Q}$. The two steps of PG can be intertwined, so that the product graph need not be materialized.
@@
In this article, we introduce OSPG, a refinement of PG that is output-sensitive in the sense that its running time depends on the size of the query output. The data complexity of OSPG is $O(|E|^{3/2}+\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$, where OUT is the query output size. OSPG achieves this complexity without knowing a priori the size of the query output. Moreover, the OSPG complexity is never higher than the complexity of PG in any case, and can be asymptotically lower in many common cases. In particular, $|E|\leq|V|^{2}$ in any graph, and $|E|\leq O(1)\cdot|V|^{2}$ in any edge-labeled graph. [^6] There can be multiple edges between the same pair of vertices, but with different labels. However, the number of labels is constant in data complexity because the query size is a constant, and we only need to consider labels that occur in the query. Hence, $|V|$ ranges from $O(|E|^{1/2})$ to $O(|E|)$, so from fully dense to very sparse graphs. Therefore, the PG complexity of $O(|V|\cdot|E|)$ ranges from $O(|E|^{3/2})$ to $O(|E|^{2})$, thus always subsuming the term $|E|^{3/2}$ in the OSPG complexity. The OSPG complexity is strictly lower when the input graph is sparse and the query output is small. This regime includes common scenarios, where the query is selective and the number of edges is far from maximum. In this regime, the improvement of OSPG over PG is due to the unnecessary time wasted by PG in the breadth-first search phase when only a few output pairs are eventually discovered.
@@ 2. Preliminaries
In the literature, given a regular path query $Q$ over an input graph $G=(V,E,\Sigma)$, it is standard to consider the product graph $P_{G,Q}$ of $G$ with an NFA $M_{Q}=(V_{Q},E_{Q},\Sigma)$ that recognizes the language $L(Q)$. The initial states of $P_{G,Q}$ are pairs $(v,q)$ where $v$ is a vertex in $G$ and $q$ is an initial state of $M_{Q}$. The accepting states of $P_{G,Q}$ are pairs $(v,q)$ where $v$ is a vertex in $G$ and $q$ is an accepting state of $M_{Q}$. The goal is to find paths from initial states to accepting states in the product graph.
Definition 2.5 ($Q$-reachability and $Q$-degree of a vertex in an edge-labeled graph).
@@ 4.1. Reduction from any RPQ $Q$ to the RPQ $ab^{*}c$
Our reduction from any RPQ $Q$ to the RPQ $ab^{*}c$ essentially shows that $ab^{*}c$ is the hardest RPQ. Let $M_{Q}=(V_{Q},E_{Q},\Sigma)$ be an NFA for $Q$, represented as an edge-labeled graph (cf. Section 2). First, we construct the product graph $\overline{G}\stackrel{{\scriptstyle\text{def}}}{{=}}G\times M_{Q}$. Let $\overline{V}$ and $\overline{E}$ be the set of vertices and edges of $\overline{G}$, respectively. Then, we construct an edge-labeled graph $G^{\prime}=(V^{\prime},E^{\prime},\Sigma^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\{a,b,c\})$ as follows: The set of vertices $V^{\prime}$ is the same as the set of vertices $\overline{V}$. The set of edges $E^{\prime}$ contains the same edges in $\overline{E}$ but with the additional label ``$b"$. Moreover, $E^{\prime}$ contains a self-loop labeled ``$a"$ for each vertex corresponding to an initial state in the NFA $M_{Q}$, and a self-loop labeled ``$c"$ for each vertex corresponding to an accepting state in $M_{Q}$. Formally:
@@
$\{((v,q),``a",(v,q))\mid\text{$v\in V$ and $q$ is an initial state in the NFA $M_{Q}$}\}\quad\cup$
$\{((v,q),``c",(v,q))\mid\text{$v\in V$ and $q$ is an accepting state in the NFA $M_{Q}$}\}$
@@
Suppose we have an edge-labeled graph $G=(V,E,\Sigma=\{d,e,f,g\})$ and consider the RPQ $Q=d^{*}(e\cdot f+g)^{*}$. We demonstrate the above reduction from $Q$ to the RPQ $ab^{*}c$ over another edge-labeled graph $G^{\prime}=(V^{\prime},E^{\prime},\Sigma^{\prime}=\{a,b,c\})$. The regular expression $Q$ can be translated into the DFA $M_{Q}=(V_{Q},E_{Q},\Sigma)$ that is depicted in Figure 1. This DFA has states $\{q_{0},q_{1},q_{2}\}$, an initial state $q_{0}$, accepting states $\{q_{0},q_{2}\}$, and the following transitions, given as labeled edges:
@@ 4.2. Solving the RPQ $ab^{*}c$
Given a vertex $x$, we define $\deg_{R}(x)$ to be the number of distinct $y$ values satisfying $R(x,y)$. In the second step of the algorithm, we partition vertices $x\in V$ based on $\deg_{R}(x)$. In particular, a vertex $x$ is called light if $\deg_{R}(x)\leq\sqrt{|E|}$ and heavy otherwise. Note that by definition of $R$ and Definition 2.5, this is equivalent to saying that a vertex $x$ is light if its $(b^{*}c)$-degree, $\deg_{b^{*}c}(x)$, is at most $\sqrt{|E|}$ and heavy otherwise. We define $R_{\ell}$ to be the set of pairs of vertices $(x,y)\in V\times V$ where $x$ is light and $y$ is $(b^{*}c)$-reachable from $x$, as shown in Eq. (3). We also define $R_{h}$ in Eq. (4) to be the set of heavy vertices $x\in V$. Given $R$, both $R_{\ell}$ and $R_{h}$ are straightforward to compute.
Finally, the algorithm evaluates the Datalog program given by Eq. (5)–(8), which defines two binary relations $Q_{\ell}$ and $Q_{h}$, and the algorithm returns their union as the final output to the RPQ. We show below that $Q_{\ell}$ contains all pairs of vertices $(v,u)$ where $u$ is $(ab^{*}c)$-reachable from $v$ through a path that only visits light vertices. In contrast, $Q_{h}$ contains all pairs of vertices $(v,u)$ where $u$ is $(ab^{*}c)$-reachable from $v$ through a path that visits at least one heavy vertex. Therefore, the algorithm computes the correct output to the RPQ $ab^{*}c$. The Datalog program contains a recursive definition for a binary relation $T$, given by Eq. (6) and (7). We evaluate $T$ using semi-naïve evaluation. In particular, at every iteration of the recursion, we only examine the new pairs of vertices $(x,z)\in T$ that were just added to $T$ in the previous iteration, and go through vertices $y$ that are $(b)$-reachable from $z$. We add the pairs $(x,y)$ to $T$ if they do not already exist. Only those newly added pairs to $T$ will be examined in the next iteration, and so on. We show in Section 5 that using this evaluation strategy, $Q_{h}$ can be computed in time $O(\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$. In contrast, $Q_{\ell}$ takes time $O(|E|^{3/2})$.
@@
4:Compute $R(X,Y)$ below while maintaining for every $X$ value, at most $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{|E|}+1$ distinct $Y$ values:
@@
7:Compute $Q_{\ell}$ and $Q_{h}$. For $Q_{h}$, use semi-naïve evaluation to compute $T$: $\triangleright$ See Section 4.2.
@@ 5. Complexity Analysis
Thanks to the reduction in Section 4.1, we only need to analyze the runtime of Algorithm 1 for the special RPQ $ab^{*}c$. We start with analyzing the runtime needed for the first step of Algorithm 1, which computes the relation $R(X,Y)$. Recall that for each vertex $x\in V$, $R$ contains at most $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{|E|}+1$ different vertices $y$ that are $(b^{*}c)$-reachable from $x$. Therefore, $|R|\leq|V|\cdot\Delta$. In order to compute $R$, we maintain, for every vertex $x$, a list $\mathsf{list}(x)$ of up to $\Delta$ different vertices $y$ that are $(b^{*}c)$-reachable from $x$, as follows:
@@
The size of $S_{h}$ cannot exceed $\min\left(\frac{\textsf{OUT}}{\sqrt{|E|}},|V|\right)$. This is because for every vertex $x\in S_{h}$, there are more than $\sqrt{|E|}$ different vertices $y$ such that the pair $(x,y)$ is in the output of the RPQ $ab^{*}c$, and the total number of such pairs $(x,y)$ is OUT by definition. By Rule (6) and then inductively by Rule (7), we have $\pi_{X}T(X,Y)\subseteq S_{h}$. By Rule (8), we have $\pi_{X}Q_{h}(X,Y)\subseteq\pi_{X}T(X,Z)\subseteq S_{h}$. For a fixed value $x\in S_{h}$ of the variable $X$, the semi-naïve evaluation of Rules (6) and (7) takes $O(|E|)$ time because it corresponds to a single-source traversal of the input graph where the source is $x$. In particular, during this single-source traversal, every vertex $z$ that is $(ab^{*})$-reachable from $x$ will be added to $T$ only once, and the outgoing edges from $z$ will be examined only once, thus leading to a total of $O(|E|)$ time. Rule (8) also takes $O(|E|)$ time for a fixed value $x$ of $X$ because the other two variables $Z$ and $Y$ form an edge, $E(Z,``c",Y)$. Hence, the total time over all possible values $x\in S_{h}$ of the variable $X$ is $O\left(\min\left(\frac{\textsf{OUT}}{\sqrt{|E|}},|V|\right)\cdot|E|\right)=O(\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$, as desired. ∎
6. When is OSPG strictly better than PG?
As mentioned in the introduction, the OSPG complexity of $O(|E|^{3/2}+\min(\textsf{OUT}\cdot\sqrt{|E|},|V|\cdot|E|))$ is never higher than the PG complexity of $O(|V|\cdot|E|)$ in any case. This is because every edge-labeled graph satisfies $|E|\leq O(1)\cdot|V|^{2}$. Hence, $|V|$ ranges from $O(|E|^{1/2})$ to $O(|E|)$, and the PG complexity ranges from $O(|E|^{3/2})$ to $O(|E|^{2})$. Moreover, the OSPG complexity is strictly lower than PG when the input graph is sparse and the query output is small. This holds specifically for $\textsf{OUT}=O((|V|\cdot\sqrt{|E|})^{1-\alpha})$ and $|E|=O(|V|^{2-\beta})$ for any $\alpha\in(0,1]$ and $\beta\in(0,2)$. This regime includes common scenarios, where the query is selective and the number of edges is far from maximum. In this regime, the improvement of OSPG over PG is due to the unnecessary time used by PG in the breadth-first search phase when only a few output pairs are eventually discovered. We next exemplify this gap between OSPG and PG in two cases.